OpenStudy (anonymous):

Related Rates Problem A man 6ft tall walks at the rate of 4 ft/sec toward a street light that is 16ft above the ground. At what rate is the length of his shadow changing when he is 7ft from the base of the light?

OpenStudy (noelgreco):

First, draw a picture and label similar triangle parts.

OpenStudy (dumbcow):

 set up proportion \[\frac{s}{6} = \frac{x+s}{16}\] solving for "s" \[s = \frac{3}{5} x\] chain rule says \[\frac{ds}{dt} = \frac{ds}{dx}*\frac{dx}{dt}\] dx/dt is given as -4 since you are walking towards light, x is decreasing \[\rightarrow \frac{ds}{dx} = \frac{3}{5}\] therefore \[\frac{ds}{dt} = \frac{3}{5}*(-4) = -\frac{12}{5}\]